OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of . (3 points) y − 6 = (x + 3) y = x − 6 y + 3 = (x − 6) x – 2y = 12

OpenStudy (anonymous):

A slope of 1/2 *

OpenStudy (anonymous):

I need to show work too guys.. I am sooo terrible at math i am sorry D;

OpenStudy (anonymous):

what math class is this

OpenStudy (anonymous):

Its algebra 1

OpenStudy (anonymous):

what unit

OpenStudy (anonymous):

4 i think. I just need the answer ehhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

\[\frac{y-y_1}{x-x_1} = m \] , where m is the slope. here, you have y1 = -3 , x1 = 6, and m = 1/2 .. Can you solve further?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

just substitute the vaues of y1,x1 and m in the equation! That's pretty simple, no?